There were 115 people who attended the eagle basketball game. Tickets for students sold for $4 each and tickets for non-students sold for $7 each. If the total ticket sales were $553, how many students attended the game?
Answer by checkley77(12844)

s+ns=115
s=115-ns
4s+7ns=553
4(115-ns)+7ns=553
460-4ns+7ns=553
3ns=553-460
3ns=93
ns=93/3
ns=31 non tickets sold.
115-31=84 student tickets sold.
Proof:
4*84+7*31=553
336+217=553
553=553